5. The body, weighing 200 kg, falls freely from a height of 1000 m. Find its kinetic energy and potential 12 seconds after the start of free fall.


1
Expert's answer
2021-05-18T11:11:14-0400

In 12 seconds, the body will achieve a speed of


"v=gt=117.6\text{ m/s}."

Meanwhile it moved down for


"h=\frac{gt^2}{2}=705.6\text{ m}."

It is


"H=y-h=1000-705.6=294.4\text{ m}."

The kinetic and potential energy:


"E_K=\frac12mv^2=1.383·10^6\text{ J}.\\\space\\
E_P=mgH=577000\text{ J}."